# VideoGLUE: Video General Understanding Evaluation of Foundation Models

### v1

@@ 4.1.2 Localizing actions
The videos in action recognition are trimmed, but actions could occur anywhere in a video in the wild. Hence, temporal and spatiotemporal action localization is also crucial to video understanding. Accordingly, we choose three datasets for the experiments: the action localization track of ActivityNet v1.3 (ANet) [16], Atomic Visual Actions (AVA) [21], and AVA-Kinetics (AVA-K) [31]. The last two require a model to localize (and recognize) actions in both time and space, and their underlying videos are movies and general Youtube videos, respectively.
@@ 4.2 Adaptation methods
4.2.1 Modifying FM architectures for downstream tasks
Given a $\textsc{fm}(\cdot)$, we can apply $\textsc{fm}(\cdot)$ to a video clip $C$ to extract a set of $k$ feature maps $\{F\}^{k}=\textsc{fm}(C),F\in\mathbb{R}^{n\times h\times w\times c}$, where $k$ is the number of endpoint layers from a FM, and $n,h,w,c$ are respectively a feature map’s length, height, width, and number of channels.
For video classification tasks, we cast a feature map $F$ as $n\times h\times w$ tokens and aggregate them into a global representation using a learnable query token $\tau$ and lightweight cross-attention layers [15]. For spatiotemporal action localization, following the standard practice [19, 48], we first detect humans on key-frames using a human detector [44], producing a set of human bounding boxes $B$. We then apply the RoI pooling operation [25] that takes both the feature map $F$ and box coordinates $B$ as inputs and outputs one feature vector per box as the query token, $\tau=\textsc{RoIPool}(F,B)$, followed by the same cross-attention layers as in video classification. For both groups of tasks, we stack a linear classifier on top of the task token’s last-layer encoding for final classification:
@@ 4.2.1 Modifying FM architectures for downstream tasks
For temporal action localization, we first perform feature extraction in a sliding window manner, resulting in a sequence of globally average pooled features $\{\textsc{AvgPool}(F_{1}),\cdots,\textsc{AvgPool}(F_{t})\}$ for each video. Following a popular choice of prior works [2, 27, 36], we employ G-TAD [57] as our task head for predicting the action category and its start and end timestamps.
@@
Figure 2: We study four adaptation methods to apply a foundation model (FM) to video understanding downstream tasks: (a) end-to-end finetuning, (b) frozen backbone evaluation, (c) frozen features with multi-layer attention pooler (MLAP), and (d) a low-rank adapter.
4.2.2 Adapting the modified FMs’ weights for downstream tasks
@@
Adapting the modified FMs to a downstream task is to tune their weights. Then, we immediately have two basic adaptation strategies: 1) full finetuning to update all weights in the original FM plus the task head and 2) freezing FM weights and only updating newly added weights. The choice of the adaptation methods depends on specific application scenarios such as computation and memory constraints. We argue that an ideal FM should perform well across various adaptation methods to support the breadth of use cases.
@@
End-to-end finetuning. End-to-end finetuning is the most common FM evaluation method for videos [1, 18, 48, 55], but it requires the deployment of a separate and possibly expensive FM for each downstream task. When finetuning all weights in the modified FMs, we limit cross-attention to a single transformer layer with 12 heads and hidden size 768. We vary learning rates and weight decays for each experiment to ensure every FM is configured to its best setup. Figure 2(a) illustrates this end-to-end finetuning.
@@
Frozen FM. Linear probing and cross-attention based pooling over frozen FM features are routinely used to test the strength of the FM representation [48, 59, 47, 22, 35]. In practice, adapting task-specific heads with a frozen FM allows us to deploy the same FM for multiple tasks. If we use light-weight heads over the FM features, then a single FM inference can serve multiple tasks efficiently in terms of both compute and memory. To this end, we examine two variations with a frozen FM, one with a single cross-attention layer and the other with multiple layers. The first results in exactly the same model architectures as in end-to-end finetuning (Figure 2(b)), and the second allows us to leverage an FM’s hierarchical features beyond its last endpoint layer (Figure 2(c)). First, the frozen features are extracted from the last $k$ layers, $F_{N-k+1}$, $F_{N-k+2}$, …, $F_{N}$. Then, attention pooling is applied between a learnable token $\tau$ and the features $F_{N-k+1}$ using multi-head cross-attention (MHCA). The output of this layer serves as the query token for the next round of attention pooling with the features $F_{N-k+2}$. This process is repeated for $k$ rounds:
$\tau_{N-k+1}$ $=\textsc{MLP}(\textsc{MHCA}(\tau,F_{N-k+1}))$ (2)
$\tau_{N-k+2}$ $=\textsc{MLP}(\textsc{MHCA}(\tau_{N-k+1},F_{N-k+2}))$
$\tau_{N}$ $=\textsc{MLP}(\textsc{MHCA}(\tau_{N-1},F_{N}))$
@@
Frozen FM with a low-rank adapter. Finally, we explore a frozen FM beyond the last $k$ layers using a low-rank adapter [23], which is a bottleneck architecture that projects a feature tensor into a low-dimensional space and then up-samples to the original space. The bottleneck space’s dimension is 64 in our experiments. Inserting a few adapter layers with trainable weights $\{w\}$ into the pretrained FM while keeping all FM’s weights frozen, the feature adapter is more parameter-efficient than end-to-end finetuning the whole network while achieving better performance than simply adding a task head to the frozen FM. Essentially, the adapter leads to a new $\widetilde{\textsc{FM}}$ with some trainable weights $\{w\}$: $\tilde{F}=\widetilde{\textsc{FM}}(C,\{w\}),$ such that the output feature maps remain the same in shape as the original FM’s output (Figure 2(d)). Hence, different pooling schemes and task heads aforementioned could be applied to the extracted feature map $\tilde{F}$. For simplicity, we still choose the single-layer cross-attention as the default task head due to its computation efficiency and performance.
@@ 5.1 End-to-end finetuning
Table 3 shows the end-to-end finetuning results of six FMs on eight datasets. We split the FMs into two groups based on their input modalities at the time of pretraining: CoCa, CLIP, and FLAVA are image-native FMs, and VideoMAE, VATT, and InternVideo are video-native. The datasets span spatiotemporal action localization (STAL), video classification (VC), and temporal action localization (TAL). Note that we freeze FM weights in TAL because otherwise its full finetuning consumes excessive memory and computation. We draw the following observations from Table 3.
@@
Table 3: Evaluating FMs when adapted to video understanding tasks using end-to-end finetuning. We report the Top-1 accuracy on K400, MiT, D48 and SSv2, MAP on Charades and ANet, and mAP@IOU0.5 on AVA and AVA-K.
STAL TAL VC (A) VC (M) VC (ML)
Method AVA AVA-K ANet K400 MiT D48 SSv2 Charades AVG
CoCa [59] 27.7 31.0 -- 82.6 43.6 79.6 $66.8$ $55.0$ $55.2$
CLIP [41] $27.1$ $28.9$- $81.0$ $39.0$ $75.7$ $46.6$ $54.3$ $52.8$
FLAVA [47] $22.0$ $25.6$- $79.1$ $38.3$ $72.0$ $61.1$ $48.6$ $49.4$
VideoMAE [18] $23.5$ $26.2$- $78.7$ $36.1$ $75.5$ $65.5$ $51.4$ $51.0$
InternVideo [55] $27.2$ $29.8$- $80.1$ $35.9$ $75.8$ 67.0 $52.2$ $52.5$
VATT [1] $27.0$ $28.4$- $77.1$ $34.8$ $77.6$ $65.1$ 55.7 $52.7$
Task-specialized $42.3$ $38.9$ $37.5$ $88.6$ $42.7$ $88.9$ $68.7$ $63.2$-
[42] [42] [54] [40] [33] [58] [17] [29]
@@
All six FMs underperform task-specialized models on the video tasks. Table 3’s last row collects the state-of-the-art results on the eight datasets, each obtained by a task-specialized model with comparable architecture or size to ours in the prior work. All six FMs significantly underform the task-specialized models on the video tasks, indicating the lack of strong video-focused FMs. This observation is in sharp contrast to what FMs have achieved on natural language [39, 3] and image understanding [11].
Video-native FMs outperform image-native FMs on SSv2, Charades, and ANet which require a model to reason along the time dimension: SSv2 actions are motion-rich, Charades has multiple actions per video, and ANet is about temporal action localization. These results strut the advantages of video-native FMs over image-native ones and, hopefully, prompt more efforts dedicating to the research of video-native FMs.
CoCa performs the best among image-native FMs on the video tasks. It actually gives rise to the highest accuracy on all datasets except SSv2, Charades, and ANet probably because CoCa, pretrained using image-text pairs, does not capture sufficient motion signals required for understanding SSv2, and it cannot handle Charades and ANet’s complex, multiple actions per video.
5.2 Frozen FMs
@@
STAL TAL VC (A) VC (M) VC (ML)
Method AVA AVA-K ANet K400 MiT D48 SSv2 Charades AVG
CoCa [59] 23.3 $24.7$ 33.0 $73.1$ $32.0$ $34.1$ $41.5$ $8.8$ $30.8$
CLIP [41] $21.1$ 25.9 32.7 75.2 32.6 $44.1$ $41.0$ $11.2$ $32.8$
FLAVA [47] $18.8$ $21.5$ 32.2 $71.3$ $29.7$ $45.9$ $40.6$ $12.6$ $31.6$
VideoMAE [18] $16.0$ $19.9$ 33.0 $65.1$ $23.0$ 59.5 $53.9$ $11.3$ $32.5$
InternVideo [55] $13.4$ $15.7$ 33.3 $69.3$ $26.3$ $55.6$ 58.2 $13.0$ $33.1$
VATT [1] $20.3$ $22.2$ 35.3 $75.1$ $32.1$ $49.7$ $57.8$ 33.3 $40.5$
@@
STAL TAL VC (A) VC (M) VC (ML)
Method AVA AVA-K ANet K400 MiT D48 SSv2 Charades AVG
CoCa [59] $24.4$ $27.0$ $33.3$ $74.2$ $37.2$ $48.4$ $45.9$ $19.6$ $45.9$
CLIP [41] 27.7 29.6 $33.9$ 77.1 39.0 $55.8$ $50.1$ $41.5$ $46.2$
FLAVA [47] $21.3$ $23.2$ $32.4$ $71.5$ $34.5$ $58.5$ $43.1$ $38.2$ $41.7$
VideoMAE [18] $19.6$ $22.1$ $33.4$ $71.7$ $32.2$ $69.6$ $57.4$ $35.9$ $43.2$
InternVideo [55] $15.9$ $17.7$ $33.6$ $73.7$ $34.7$ 71.9 60.3 $40.5$ $44.8$
VATT [1] $22.9$ $24.1$ $35.0$ $75.1$ $35.6$ $60.1$ $58.7$ 58.2 $46.9$
@@
STAL VC (A) VC (M) VC (ML)
Method AVA AVA-K K400 MiT D48 SSv2 Charades AVG
CoCa [59] 26.6 28.7 80.9 41.4 $67.1$ $56.1$ $45.8$ $49.0$
CLIP [41] $24.5$ $28.0$ $80.2$ $39.7$ 77.2 $56.0$ $44.2$ $49.3$
FLAVA [47] $17.9$ $23.8$ $74.7$ $34.1$ $68.4$ $52.1$ $40.8$ $44.1$
VideoMAE [18] $16.6$ $23.3$ $73.6$ $30.6$ $76.0$ $61.4$ $43.0$ $45.9$
InternVideo [55] $19.2$ $25.5$ $75.5$ $31.3$ $73.6$ 63.9 $46.2$ $47.7$
VATT [1] $22.3$ $25.8$ $75.0$ $36.5$ $68.9$ $63.5$ 53.5 $49.9$
@@
CLIP generally performs the best among image-native frozen FMs (Tables 4 and 5), but CoCa catches up thanks to the low-rank adapter (Table 6). It is worth noting that this ranking of image-native frozen FMs differs from the ranking of image-native FMs in end-to-end finetuning. It seems that CLIP’s endpoint features are more amendable to the video tasks than CoCa, but CoCa as a whole adapts better to video under both finetuning and the adapter. Hence, it is crucial to consider adaptation methods as an organic part of the evaluation of FMs to supply them various paths to demonstrate their capabilities.
Figure 3: FMs are equipped with different adaptation methods. Left: For each adaptation method, we plot FMs’ averaged scores across all video tasks vs. trainable FLOPs in a log scale. Right: We plot the overall VideoGLUE score (VGS) per FM.
@@
Video-native FMs are better than image-native FMs in understanding motion-rich SSv2 and D48, Charades that contain multiple actions per video, and ANet for temporal action localization. This observation is about the same as the one under end-to-end finetuning. The image-native FMs is mainly superior on appearance-rich video datasets, where high-quality spatial perceptual features are the key. We conjecture that the vast image data empowering image-native FMs is more diverse in appearance than videos used to pretrain video-native FMs.
@@
Given frozen FMs, the low-rank adapter outperforms cross-attention layers, and multiple layers of cross-attention is better than a single cross-attention layer. Many works [10, 22] have shown features from different layers of a vision transformer have different attention maps. Hence, it is potentially beneficial to have an adaptation method to leverage multiple layers of a frozen FM. Table 5 reports the results with four cross-attention layers, whose average score per model (across different columns) is higher than that with a single cross-attention layer (Table 4) by $18\%$ to $40\%$. The low-rank adapter (Table 6) further improves upon the cross-attention results partially because it explores all layers of a frozen FM.
@@
On average, image-native FMs outperform video-native FMs under end-to-end finetuning and the adapter, but it becomes the inverse in the other two adaptation methods. The adapter experiment paired with end-to-end finetuning experiment reveal the fact that existing image-based FMs could be more easily adapted to video tasks when we could adjust the feature space of FMs, possibly caused by the large-scale higher quality image(-text) pretraining datasets. On the other hand, frozen feature experiments discussed above present us the inverse picture where video-based FM performs better. The seemingly paradox encourages more future research on bridging the gap on video-based pretraining with high-quality data and more effective modeling.
5.3 VideoGLUE score: Ranking FMs by their efficacies and efficiencies on video tasks
In this section, we consolidate our studies of the FMs with different adaptation methods on a broad range of video tasks by considering their adaptation efficacies and efficiencies. Adaptation methods with different numbers of trainable weights lead to incompatible comparisons. Motivated by this, we propose a scalar measure, called VideoGLUE score (VGS), to capture an FM’s overall adaptation performance on our video understanding tasks. While the VideoGLUE score may not be a perfect metric, it condenses multiple aspects of comparison into a scalar value, enabling a simplified comparison of FMs.
Taking the adaptation efficiency into account, we propose to use the trainable FLOPs to normalize an adapted FM’s average score $s$ over all tasks. The trainable FLOPs are better than tunable weights because they allow our VGS to reflect both the model architecture’s freedom and the input data’s impact (e.g., sequence length) on downstream tasks. Formally, denoting by $\mathcal{S}_{i}$ an FM’s average score over our video tasks under the $i$-th adaptation method and by $F_{i}$ the corresponding trainable FLOPs (in billion), we calculate the FM’s VGS by
[…]
where $N=4$ is the number of adaptation methods, and $w_{i}\in[0,1]$ weigh score $\mathcal{S}_{i}$ according to the trainable FLOPs $F_{i}$.

### v3

@@ 3.2.2 Localizing actions
The videos in action recognition are trimmed, but actions could occur anywhere in a video in the wild. Hence, temporal and spatiotemporal action localization is also crucial to video understanding. Accordingly, we choose three datasets for the experiments: the action localization track of ActivityNet v1.3 (ANet) (Fabian Caba Heilbron & Niebles, 2015), Atomic Visual Actions (AVA) (Gu et al., 2018), and AVA-Kinetics (AVA-K) (Li et al., 2020). The last two require a model to localize and recognize actions in both time and space, and their underlying videos are movies and general YouTube videos, respectively.
@@ 3.3 Adaptation methods
3.3.1 Modifying foundation model architectures for downstream tasks
Given an $\textsc{fm}(\cdot)$, we can apply $\textsc{fm}(\cdot)$ to a video clip $C$ to extract a set of $k$ feature maps $\{F\}^{k}=\textsc{fm}(C),F\in\mathbb{R}^{n\times h\times w\times c}$, where $k$ is the number of endpoint layers from an FM, and $n,h,w,c$ are respectively a feature map’s length, height, width, and number of channels.
For video classification tasks, we cast a feature map $F$ as $n\times h\times w$ tokens and aggregate them into a global representation using a learnable query token $\tau$ and lightweight cross-attention layers (Dosovitskiy et al., 2020). For spatiotemporal action localization, following the standard practice (Feichtenhofer et al., 2019; Tong et al., 2022), we first detect humans on key-frames using a human detector (Ren et al., 2015), producing a set of human bounding boxes $B$. We then apply the RoI pooling operation (Jaderberg et al., 2015) that takes both the feature map $F$ and box coordinates $B$ as inputs and outputs one feature vector per box as the query token, $\tau=\textsc{RoIPool}(F,B)$, followed by the same cross-attention layers as in video classification. For both groups of tasks, we stack a linear classifier on top of the task token’s last-layer encoding for final classification:
@@ 3.3.1 Modifying foundation model architectures for downstream tasks
For temporal action localization, we first perform feature extraction in a sliding window manner, resulting in a sequence of globally average pooled features $\{\textsc{AvgPool}(F_{1}),\cdots,\textsc{AvgPool}(F_{t})\}$ for each video. Following a popular choice of prior works (Alwassel et al., 2021; Ju et al., 2022; Liu et al., 2022), we employ G-TAD (Xu et al., 2020) as our task head for predicting the action category and its start and end timestamps.
@@
Figure 2: We study four adaptation methods to apply a foundation model (FM) to video understanding downstream tasks: (a) end-to-end finetuning, (b) frozen backbone, (c) frozen backbone with multi-layer attention pooler (MLAP), and (d) a low-rank adapter.
3.3.2 Adapting modified foundation model to downstream tasks
@@
Adapting the modified FMs to a downstream task is to tune their weights. Then, we immediately have two basic adaptation strategies: (1) full finetuning to update all weights in the original FM plus the task head and (2) freezing FM weights and only updating newly added weights. The choice of the adaptation methods depends on specific application scenarios such as computation and memory constraints. We argue that an ideal FM should perform well across various adaptation methods to support the breadth of use cases.
@@
End-to-end finetuning. End-to-end finetuning is the most common FM evaluation method for videos (Akbari et al., 2021; Feichtenhofer et al., 2022; Tong et al., 2022; Wang et al., 2022b), but it requires the deployment of a separate and possibly expensive FM for each downstream task. When finetuning all weights in the modified FMs, we limit cross-attention to a single transformer layer with 12 heads and hidden size 768. We vary learning rates and weight decays for each experiment to ensure every FM is configured to its best setup. Figure 2(a) illustrates this end-to-end finetuning.
@@
Freezing foundation model weights. Linear probing and cross-attention based pooling over frozen FM features are routinely used to test the strength of the FM representation (Tong et al., 2022; Yu et al., 2022; Singh et al., 2022; He et al., 2022; Lin et al., 2022). In practice, adapting task-specific heads with a frozen FM allows us to deploy the same FM for multiple tasks. If we use light-weight heads over the FM features, then a single FM inference can serve multiple tasks efficiently in terms of both compute and memory. To this end, we examine two variations with a frozen FM, one with a single cross-attention layer and the other with multiple layers. The first results in exactly the same model architectures as in end-to-end finetuning (Figure 2(b)), and the second allows us to leverage an FM’s hierarchical features beyond its last endpoint layer (Figure 2(c)). First, the frozen features are extracted from the last $k$ layers, $F_{N-k+1}$, $F_{N-k+2}$, …, $F_{N}$. Then, attention pooling is applied between a learnable token $\tau$ and the features $F_{N-k+1}$ using multi-head cross-attention (MHCA). The output of this layer serves as the query token for the next round of attention pooling with the features $F_{N-k+2}$. This process is repeated for $k$ rounds:
$\tau_{N-k+1}$ $=\textsc{MLP}(\textsc{MHCA}(\tau,F_{N-k+1})),$ (2)
$\tau_{N-k+2}$ $=\textsc{MLP}(\textsc{MHCA}(\tau_{N-k+1},F_{N-k+2})),$
$\tau_{N}$ $=\textsc{MLP}(\textsc{MHCA}(\tau_{N-1},F_{N})),$
@@
Freezing foundation model weights with low-rank adaptation. Finally, we explore a frozen FM beyond the last $k$ layers using a low-rank adapter (Hu et al., 2021), which is a bottleneck architecture that projects a feature tensor into a low-dimensional space and then up-samples to the original space. The bottleneck space’s dimension is 64 in our experiments. Through inserting a few adapter layers with trainable weights $\{w\}$ into the pretrained FM while keeping all FM’s weights frozen, the feature adapter is more parameter-efficient than end-to-end finetuning the whole network while achieving better performance than simply adding a task head to the frozen FM. Essentially, the adapter leads to a new FM with some trainable weights $\{w\}$: $\tilde{F}=\widetilde{\textsc{FM}}(C,\{w\}),$ such that the output feature maps remain the same in shape as the original FM’s output (Figure 2(d)). Hence, different pooling schemes and task heads aforementioned could be applied to the extracted feature map $\tilde{F}$. For simplicity, we still choose the single-layer cross-attention as the default task head due to its computation efficiency and performance.
@@ 4.1 End-to-end finetuning
Table 3 shows the end-to-end finetuning results of six FMs on eight datasets. We split the FMs into two groups based on their input modalities at the time of pretraining: CLIP, FLAVA, CoCa, and DINOv2 are image-native FMs, while VATT, VideoMAE, and InternVideo are video-native. The datasets span video classification (VC) and spatiotemporal action localization (STAL). Note that it is infeasible to end-to-end fine-tune or LoRA fine-tune the vision encoder on TAL task, because the videos in ANet are typically long (up to 30 minutes). We follow the common practice of pre-computing visual features in a sliding window manner offline, and training a temporal detection network on top of the visual features (Wang et al., 2021; Zhang et al., 2022). We will report TAL results in the next section. We draw the following observations from Table 3.
@@
Table 3: Evaluating FMs when adapted to video understanding tasks using end-to-end finetuning. We report the Top-1 accuracy on K400, MiT, SSv2 and D48, MAP on Charades and ANet, and mAP@IOU0.5 on AVA and AVA-K.
Models VC-A) VC-M VC-ML TAL STAL Avg.
K400 MiT SSv2 D48 Charades ANet AVA AVA-K
CLIP $81.0$ $39.0$ $46.6$ $75.7$ $54.3$- $27.1$ $28.9$ $52.8$
FLAVA $79.1$ $38.3$ $61.1$ $72.0$ $48.6$- $22.0$ $25.6$ $49.4$
CoCa 82.6 43.6 $66.8$ 79.6 $55.0$- 27.7 31.0 $55.2$
DINOv2 $80.3$ $38.7$ $60.9$ $70.2$ $52.4$- $24.6$ $27.0$ $50.5$
VATT $77.1$ $34.8$ $65.1$ $77.6$ 55.7 -- $27.0$ $28.4$ $52.7$
VideoMAE $78.7$ $36.1$ $65.5$ $75.5$ $51.4$- $23.5$ $26.2$ $51.0$
InternVideo $80.1$ $35.9$ 67.0 $75.8$ $52.2$- $27.2$ $29.8$ $52.5$
Task- $88.6$ $42.7$ $68.7$ $88.9$ $63.2$ $37.5$ $42.3$ $38.9$-
specialized (TubeViT) (UniformerV2) (MViT) (AIM) (MoViNet) (PRN) (RAFT) (RAFT)
@@
FMs underperform task-specialized models on video tasks in general. Table 3’s last row collects the state-of-the-art results on the eight datasets, each obtained by a task-specialized model with comparable architecture or size to ours in the prior work. Specifically, those task-specialized models are RAFT (Rajasegaran et al., 2023), PRN (Wang et al., 2021), TubeViT (Piergiovanni et al., 2023), UniformerV2 (Li et al., 2022b), AIM (Yang et al., 2023), MViT (Fan et al., 2021), and MoViNet (Kondratyuk et al., 2021), respectively. All seven FMs underperform the task-specialized models on all video tasks except on Moments in Time at the comparable model scale, indicating the lack of strong video-focused FMs. This observation is in sharp contrast to what FMs have achieved on natural language (OpenAI, 2022; Anil et al., 2023) and image understanding (Chen et al., 2022).
CoCa performs the best among image-native FMs on the video tasks. It actually gives rise to the highest accuracy on all datasets, with slightly inferior performance on SSv2 and Charades. This shows strong generalization capability of the CoCa model, regardless it is an image-based model with image-only pre-training data. However, in the latter session, we will reveal that under light-weight and parameter efficient adaptation scenarios, the same model may perform inferior on many video understanding tasks, especially on SSv2 (Tables 45, and 6), Charades (Tables 45, and 6), and ANet (Tables 4, and 5), which require complex motion or multiple actions understanding per video. These are in contrast that CoCa achieves the best general performance in end-to-end fine-tuning (Table 3), highlighting the importance of considering adaptation methods on FMs benchmarking.
4.2 Freezing foundation models
@@
Models VC-A VC-M VC-ML TAL STAL Avg.
K400 MiT SSv2 D48 Charades ANet AVA AVA-K
CLIP 75.2 32.6 $41.0$ $44.1$ $11.2$ $32.7$ $21.1$ $25.9$ $32.8$
FLAVA $71.3$ $29.7$ $40.6$ $45.9$ $12.6$ $32.2$ $18.8$ $21.5$ $31.7$
CoCa $73.1$ $32.0$ $41.5$ $34.1$ $8.8$ $33.0$ $23.3$ $24.7$ $31.2$
DINOv2 $72.3$ $29.0$ $40.0$ $40.4$ $25.8$ $32.6$ 24.6 26.9 $35.0$
VATT $75.1$ $32.1$ $57.8$ $49.7$ 33.3 35.3 $20.3$ $22.2$ $39.1$
VideoMAE $65.1$ $23.0$ $53.9$ 59.5 $11.3$ $33.0$ $16.0$ $19.9$ $32.6$
InternVideo $69.3$ $26.3$ 58.2 $55.6$ $13.0$ $33.3$ $13.4$ $15.7$ $33.1$
@@
Models VC-A VC-M VC-ML TAL STAL Avg.
K400 MiT SSv2 D48 Charades ANet AVA AVA-K
CLIP 77.1 39.0 $50.1$ $55.8$ $41.5$ $33.9$ 27.7 29.6 $43.3$
FLAVA $71.5$ $34.5$ $43.1$ $58.5$ $38.2$ $32.4$ $21.3$ $23.2$ $39.3$
CoCa $74.2$ $37.2$ $45.9$ $48.4$ $19.6$ $33.3$ $24.4$ $27.0$ $36.3$
DINOv2 $75.4$ $36.0$ $46.3$ $51.9$ $47.8$ $33.6$ $25.4$ $27.0$ $42.5$
VATT $75.1$ $35.6$ $58.7$ $60.1$ 58.2 35.0 $22.9$ $24.1$ $46.3$
VideoMAE $71.7$ $32.2$ $57.4$ $69.6$ $35.9$ $33.4$ $19.6$ $22.1$ $40.9$
InternVideo $73.7$ $34.7$ 60.3 71.9 $40.5$ $33.6$ $15.9$ $17.7$ $42.2$
@@
Models VC-A VC-M VC-ML TAL STAL Avg.
K400 MiT SSv2 D48 Charades ANet AVA AVA-K
CLIP $80.2$ $39.7$ $56.0$ 77.2 $44.2$- $24.5$ $28.0$ $49.3$
FLAVA $74.7$ $34.1$ $52.1$ $68.4$ $40.8$- $17.9$ $23.8$ $44.1$
CoCa 80.9 41.4 $56.1$ $67.1$ $45.8$- 26.6 28.7 $49.0$
DINOv2 $77.7$ $36.1$ $59.0$ $76.6$ $38.0$- $22.5$ $27.9$ $47.0$
VATT $75.0$ $36.5$ $63.5$ $68.9$ 53.5 -- $22.3$ $25.8$ $49.9$
VideoMAE $73.6$ $30.6$ $61.4$ $76.0$ $43.0$- $16.6$ $23.3$ $45.9$
InternVideo $75.5$ $31.3$ 63.9 $73.6$ $46.2$- $19.2$ $25.5$ $47.7$
@@
Generally speaking, DINOv2 performs the best in the frozen feature pooler evaluation (Tables 4), CLIP performs the best with multi-head attention pooler among image-native frozen FMs (Tables 5), but CoCa catches up thanks to the low-rank adapter (Table 6). It is worth noting that this ranking of image-native frozen FMs differs from the ranking of image-native FMs in end-to-end finetuning. It seems that DINOv2 and CLIP’s frozen features are more amendable to the video tasks than CoCa, but CoCa as a whole adapts better to video under both finetuning and the adapter. Hence, it is crucial to consider adaptation methods as an organic part of the evaluation of FMs to supply them various paths to demonstrate their capabilities.
Figure 3: (a) We measures the training (red diamond) and inference (blue square) cost of different adaptation methods in terms of number of trainable parameters and inference FLOPs, respectively. (b) We report VideoGLUE Score that combines a FM’s performance weighted by its training costs with different adaptation methods for all the image-native (red circle) and video-native (blue pentagon) models.
@@
Video-native FMs are better than image-native FMs in understanding motion-rich SSv2 and D48, Charades that contain multiple actions per video, and ANet for temporal action localization. This observation is about the same as the one under end-to-end finetuning. The image-native FMs are mainly superior on appearance-rich video datasets, where high-quality spatial perceptual features are the key. We conjecture that the vast image data empowering image-native FMs is more diverse in appearance than videos used to pretrain video-native FMs.
@@
Given frozen FMs, the low-rank adapter outperforms cross-attention layers, and multiple layers of cross-attention is better than a single cross-attention layer. Many works (Caron et al., 2021; He et al., 2022) have shown features from different layers of a vision transformer have different attention maps. Hence, it is potentially beneficial to have an adaptation method to leverage multiple layers of a frozen FM. Table 5 reports the results with four cross-attention layers, whose average score per model (across different columns) is higher than that with a single cross-attention layer (Table 4) by $18\%$ to $40\%$. The low-rank adapter (Table 6) further improves upon the cross-attention results partially because it explores all layers of a frozen FM.
@@
On average, image-native FMs outperform video-native FMs under end-to-end finetuning and the adapter, but it becomes the inverse in the other two adaptation methods. The adapter experiment paired with end-to-end finetuning experiment reveal the fact that existing image-based FMs could be more easily adapted to video tasks when we could adjust the feature space of FMs, possibly caused by the large-scale higher quality image(-text) pretraining datasets. On the other hand, frozen feature experiments discussed above present us the inverse picture where video-based FMs perform better. The seemingly paradox encourages more future research on bridging the gap on video-based pretraining with high-quality data, more effective modeling and better design on video benchmarks.
4.3 Profiling foundation models for video understanding
In this section, we consolidate our studies of the FMs with different adaptation methods and video tasks, focusing on their overall efficacy and efficiency. Specifically, we use trainable parameters and inference FLOPs to approximately represent the training and inference costs of an FM. Since all FMs in our evaluation are ViT-B and we align the same number of input tokens for each task. The models have almost the same cost in one adaptation method. The left of Figure 3 shows the cost values for each adaptation method. Note that an FM with LoRA adaptor tuning could have high inference cost despite lower training/adaptation costs than end-to-end fine-tuning. While the figure provides a holistic view of an FM from multiple dimensions, one might be interested in a ranking among the FMs in terms of their video understanding capabilities. To this end, we summarize the multi-dimensional comparisons across different datasets, adaptation methods, and costs using a simplified scalar measure, termed VideoGLUE Score (VGS), to probe an FM’s general video understanding capability.
We use the cost values to normalize an adapted FM’s average performance $s$ over all tasks. Formally, denoting by $\mathcal{S}_{i}$ an FM’s average performance score over our video tasks under the $i$-th adaptation method and by $C_{i}^{k}$ the corresponding cost value under the $k$-th developmental scenario, we calculate the FM’s VGSk by
$\textit{VGS}^{k}=\sum_{i=1}^{N}w_{i}^{k}\mathcal{S}_{i},\;\text{where}\;w_{i}^{k}=\frac{\mathcal{A}_{i}^{k}}{\sum_{j=1}^{N}\mathcal{A}_{j}^{k}}\;\text{and}\;\mathcal{A}_{i}^{k}=\frac{1}{\log_{10}{C_{i}^{k}}},$ (3)
where $N=4$ is the number of adaptation methods, and $w_{i}\in[0,1]$ weighs score $\mathcal{S}_{i}$ according to the cost $C_{i}^{k}$. The final VGS is the arithmetic average on $\{\text{VGS}^{k}\}$, where $k=1,2$ corresponding to training and inference respectively.